A metal nitrate reacts with KI to give a black precipitate which on addition of excess of KI is converted into orange colour solution. The cation of the metal nitrate is
a. Hg^2+
b. Bi^3+
c. Pb^2+
d. Cu^+

Grade:12th pass

1 Answers

bharat bajaj IIT Delhi
9 years ago
Bi(NO3)3 + 3KI = BiI3 + 3KNO3
BiI3 + KI = K(BiI4) orange
b) is the answer
THanks
Bharat Bajaj
IIT Delhi
askiitians faculty